OpenStudy (anonymous):

Solve the system y = x + 2 and y = –x – 3 (2 , 4) (–1 , –2) (–2.5 , –0.5) (3.5 , –4.5)

OpenStudy (anonymous):

substitute y=x+2 to the second equation and it comes... x+2=-x-3 x+x=-3-2 2x=-5 x=-5/2 now substitute back x=-5/2 into y=x+2 and you have y=(-5/2)+2 y=-1/2 so x=-5/2 y=-1/2 or x=-2.5 and y=-0.5 so the answer is C :)

OpenStudy (anonymous):

thank you Kreshnik :)

OpenStudy (anonymous):

you're welcome.. Glad to help :)